OpenStudy (anonymous):

Determine the sum of a geometric series whose first term is 5, whose common ratio is 3/5 and whose number of terms is infinite.

OpenStudy (anonymous):

12.5

OpenStudy (anonymous):

that is what i got but did not think it was right because it was not a choice on my answers but o well it could be none of the above THANKS MATE!

OpenStudy (anonymous):

= 5 / (1- 3/5) = 5 / 2/5 = 5/1 * 5/2 = 25 / 2 = 12.5

OpenStudy (anonymous):

what about this one could you help me on this type?

OpenStudy (anonymous):

Determine the sum of a geometric series with 12 terms whose first term is 2 and whose second term is −4.

OpenStudy (anonymous):

-2730

OpenStudy (anonymous):

thanks can u help me on the other one i posted i need to check my anser this is not a multiple choice one